OpenStudy (anonymous):

A single die is rolled. Find the probability of rolling an even number or a number less than 4

OpenStudy (ranga):

Rolling a die has 6 possible outcomes. Getting an even number means getting a 2, 4, or 6. So 3 possibilities. Probability of getting an even number is: 3/6 = 1/2 contd...

OpenStudy (ranga):

Number less than 4 means: 1, 2, and 3. 3 possibilities Probability of getting a number less than 4 is: 3 / 6 = 1/2 These two events are not mutually exclusive because 2 occurs in both cases. The probability of getting a 2 is: 1/6 P(A or B) = P(A) + P(B) - P(A and B)

OpenStudy (ranga):

P(A or B) = 1/2 + 1/2 - 1/6 = 5/6

OpenStudy (anonymous):

WAY TO GO RANGA!!!! you explained it well :)

OpenStudy (ranga):

Thank you Victoria.

OpenStudy (anonymous):

haha no problem

OpenStudy (anonymous):

Wow. All that for simple answer. Sometimes I hate math. -_-